OpenStudy (anonymous):

6(4)-3(5)+(15-3) ______________________yes the line is suppose to 3^2-2 be there. Need steps.

OpenStudy (anonymous):

line is divide or ?? all in one line?

OpenStudy (anonymous):

line is divide

OpenStudy (anonymous):

sorry is 3 only

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

answer is 3

OpenStudy (anonymous):

\[(6*4 - 3*5 + 12) \div (3^{2} - 2)\] = (24 -15 + 12) / (9- 2) =(21) / (7) =3